OpenStudy (mackenzie2013):

A group of students were surveyed regarding their enjoyment of country and jazz music. 8 students enjoyed both types of music, giving a total of 32 students who enjoyed country music, and 14 who enjoyed jazz. How many total students were questioned in this survey? 54 38 46 26

OpenStudy (anonymous):

32+14-8 = 46-8 = 38

OpenStudy (mackenzie2013):

So you subtract the both (8) to thirty two and fourteen to get the answer?

OpenStudy (anonymous):

yes you do.